# Lattice isomorphic Banach lattices of polynomials

### v1

@@ Abstract
We study Díaz-Dineen’s problem for regular homogeneous vector-valued polynomials. In particular, we prove that whenever $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We also study the analogous problem for the classes of regular compact, regular weakly compact, orthogonally additive and regular nuclear polynomials.
@@ 1 Introduction and background
In [15], Díaz and Dineen posed the following question: If $X$ and $Y$ are dual isomorphic Banach spaces, that is $X^{\ast}$ and $Y^{\ast}$ are isomorphic as Banach spaces, are $\mathcal{P}(^{n}X)$ and $\mathcal{P}(^{n}Y)$ isomorphic as Banach spaces? Although it is not known if the above question has a positive answer, some partial answers were presented in [13, 15, 22]. In particular, J. C. Díaz and S. Dineen proved that the problem has a positive solution whenever $X^{\ast}$ has both the Schur property and the approximation property (see [15, Proposition 4]). In [13], F. Cabello Sánchez, J. Castillo and R. García generalized Díaz-Dineen results by proving a positive solution for this problem whenever $X$ is Arens-regular or whenever $X$ and $Y$ are stable. The same result was proven, independently, by S. Lassalle and I. Zalduendo in [22] who also addressed the same question concerning other classes polynomials. Further, in [14], D. Carando and S. Lassalle extended these results to spaces of vector-valued homogeneous polynomials.
@@
Motivated by the increasing interest in studying homogeneous polynomials on Banach lattices (see, e.g., [5, 6, 7, 8, 9, 10, 18, 24]), the following question arises naturally: Given two dual isomorphic Banach lattices $E$ and $F$, that is $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic, under which conditions $\mathcal{P}^{r}(^{n}E)$ and $\mathcal{P}^{r}(^{n}F)$ are lattice isomorphic for every $n\in\mathbb{N}$? By adapting the proofs from [14] and [22], we were able to provide a positive answer whenever $E^{*}$ and $F^{*}$ have order continuous norms. In particular, we prove that whenever $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having order continuous norm then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We observe that, in our setting, the assumption of Arens regularity is replaced with the requirement that the norm on the dual is order continuous. This replacement is justified by the fact that a positive version of Arens regularity for $E$ is equivalent to the order continuity of the norm on $E^{*}$ having order continuous norm. Furthermore, as we will demonstrate in Proposition 2.1, positive Arens regularity coincides with positive symmetric Arens regularity, highlighting a noteworthy contrast with the Banach space case.
@@
As an application, we obtain conditions so that $\mathcal{P}_{\cal K}^{r}(^{n}E;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}E;G^{*})$) and $\mathcal{P}_{\cal K}^{r}(^{n}F;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}F;G^{*})$) are isomorphic, where $\mathcal{P}_{\cal K}^{r}$ and $\mathcal{P}_{\cal W}^{r}$ denote, respectively, the classes of regular compact and regular weakly compact polynomials. We also consider the class of orthogonally additive polynomials. In Section 3, we investigate this problem for a new class of polynomials. Specifically, inspired by ideas from [4], we define the class of regular nuclear $n$-homogeneous polynomials, denoted by $\mathcal{P}_{\cal N}^{r}$, and prove that whenever $G$ is Dedekind complete with the lattice approximation property and $E^{*}$ is lattice isomorphic to $F^{*}$, then $\mathcal{P}_{\cal N}^{r}(^{n}E;G)$ and $\mathcal{P}_{\cal N}^{r}(^{n}F;G)$ are lattice isomorphic for every $n\in\mathbb{N}$.
@@
Before proceeding, we introduce some definitions and notation. Throughout this paper, $X$ and $Y$ denote Banach spaces, $E$ and $F$ denote Banach lattices, $E^{+}$, $B_{E}$ and $S_{E}$ denote, respectively, the positive cone, the closed unit ball, and the unit sphere of $E$. Given Banach spaces $X_{1},\dots,X_{n}$ and $Y$, the Banach space of all continuous $n$-linear operators $A\colon X_{1}\times\cdots\times X_{n}\to Y$ is denoted by $\mathcal{L}(X_{1},\dots,X_{n};Y)$. In the case that $F=\mathbb{K}$, we denote simply by $\mathcal{L}(X_{1},\dots,X_{n})$. For given Banach lattices $E_{1},\dots,E_{n}$ and $F$, a $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$ is said to be positive if $A(x_{1},\dots,x_{n})\geq 0$ for all $x_{1}\in E_{1}^{+},\dots,x_{n}\in E_{n}^{+}$. The difference of two positive $n$-linear operators is called a regular $n$-linear operator, and the set of all regular $n$-linear operators from $E_{1}\times\cdots\times E_{n}$ into $F$ is denoted by $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$. Whenever $F$ is Dedekind complete, $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$ is a Banach lattice with the regular norm $\|A\|_{r}=\||A|\|$, where $|A|$ denotes the absolute value of the regular $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$. The positive projective tensor product, or the Fremlin tensor product of $E_{1},\dots,E_{n}$, denoted by $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$, is the Banach lattice generated by considering the completion of the vector lattice tensor product $E_{1}\overline{\otimes}\cdots\overline{\otimes}\,E_{n}$ with respect the positive projective tensor norm:
@@
is an isometric isomorphism and a lattice hommomorphism (see [11, Proposition 3.3]).
An $n$-homogeneous polynomial between Riesz spaces $P\colon E\to F$ is positive if its associated symmetric multilinear operator $T_{P}\colon E^{n}\to F$ is positive. The difference of two positive $n$-homogeneous polynomials is called a regular homogeneous polynomial, and the set of all these polynomials is denoted by $\mathcal{P}^{r}(^{n}E,F)$. When $F$ is the scalar field we simply write $\mathcal{P}^{r}(^{n}E)$. If $E$ and $F$ are Banach lattices with $F$ Dedekind complete, then $\mathcal{P}^{r}(^{n}E,F)$ is a Banach lattice with the regular norm $\|P\|_{r}=\||P|\|$, where $|P|$ denotes the absolute value of the regular $n$-homogeneous polynomial $P\colon E\to F$. The $n$-fold positive projective symmetric tensor product of $E$ by $\widehat{\otimes}_{n,s,|\pi|}E$, which is a Banach lattice endowed with the positive projective symmetric tensor norm:
@@
where $\otimes^{n}x=x\,\otimes\stackrel{{\scriptstyle n}}{{\cdots}}\otimes\,x$ for every $x\in E$. Let $\theta_{n}\colon E\to\widehat{\otimes}_{n,s,|\pi|}E$ be the canonical $n$-homogeneous polynomial given by $\theta_{n}(x)=\otimes^{n}x$. We note that $\theta_{n}$ is a lattice homomorphism. For every $P\in\mathcal{P}^{r}(^{n}E,F)$ there exists a unique regular linear operator $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$, called the linearization of $P$, such that $P(x)=P^{\otimes}(\otimes^{n}x)$ for every $x\in E$. The operator
@@
is an isometric isomorphism and a lattice homomorphism. For each $P\in\mathcal{P}^{r}(^{n}E;F)$, the regular linear operator $\Phi^{-1}(P)\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$ is called the linearization of $P$, and it is denoted by $P^{\otimes}$ (see [11, Proposition 3.4]).
We refer the reader to [1, 26] for background on Banach lattices, to [17] for Banach space theory, to [16] for polynomials on Banach spaces, and to [11, 25] for positive/regular multilinear operators or polynomials.
@@ 2 Díaz-Dineen problem for regular polynomials
This process is repeated until we obtain $\overline{A}:=A^{(n+1)*}\colon E_{1}^{**}\times\cdots\times E_{n}^{**}\to F^{**}$, the $(n+1)$-Arens adjoint of $A$ which is called the Arens-extension of $A$ (see [8, Section 2] for more details). For the remainder of the paper, we shall call this extension the Arens extension, bearing in mind that it coincides with its Aron-Berner extension (see [3]). Whenever $X=Y$, the transpose of $A$ is the bilinear map $A^{t}\colon X\times X\to Z$ given by $A^{t}(x,y)=A(y,x)$ for all $x,y\in X$. Still under the assumption that $X=Y$, AA is said to be Arens-regular (or simply regular) if $A^{t***t}=A^{***}$. It is important to note that we have two different concepts of regularity for multilinear operators, one in the lattice sense presented in the Introduction and the one in the extension sense. It follows from [29, Theorem 2.2] that a bilinear form $A\colon X\times X\to\mathbb{R}$ is Arens-regular if and only if the bounded linear operator $S_{A}\colon X\to X^{*}$ given by $S_{A}(x)(y)=A(x,y)$ is weakly compact.
Recall that a Banach space $X$ is called Arens-regular if all linear operators $X\to X^{\ast}$ are weakly compact. To our knowledge, the first appearance of a positive version of Arens-regularity was in [12], when G. Buskes and R. Page gave necessary and sufficient conditions so that every positive bilinear form $E\times E\to\mathbb{R}$ is Arens-regular. The next result summarizes what we have about this notion.
@@
Since $B$ is positive, we may consider the positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ given by $A(x,y)=B(\Pi(x),\Pi(y))$ for all $x,y\in E$. Let us see that $A$ cannot be Arens-regular, showing that its associated operator $S_{A}\colon E\to E^{*}$ is not weakly compact: as $\{S_{B}(e_{2n}):n\in\mathbb{N}\}$ is not a weakly precompact subset of $\ell_{1}^{*}=\ell_{\infty}$ (see [3, p. 83]), there exists $\varphi\in\ell_{\infty}$ such that $(\varphi(S_{B}(e_{2n})))_{n}$ does not have any Cauchy subsequence. If $x_{n}\in E$ is such that $\Pi(x_{n})=e_{2n}$, then $(\varphi(S_{B}(\Pi(x_{n}))))_{n}$ does not have any Cauchy subsequence. By extending $\varphi\colon\ell_{1}\to\mathbb{R}$ to $E$ in such a way that it vanishes on the complement, we obtain that $\{\Pi^{*}\circ S_{B}\circ\Pi(x_{n}):n\in\mathbb{N}\}$ is not weakly precompact in $E^{*}$, and since $S_{A}=\Pi^{*}\circ S_{B}\circ\Pi$, we conclude that $S_{A}$ is not weakly compact. ∎
The fourth condition in Proposition 2.1 induces a positive version of Arens-regularity. However, since this is equivalent to a Banach lattice whose dual has order continuous norm, we will not introduce a new terminology. It is noteworthy mentioning that item (5) in Proposition 2.1 shows that the notions of positive Arens-regularity and positive symmetric Arens-regularity are equivalent. This result highlights an important difference between the theory of Banach spaces and the theory of Banach lattices. In [23], D. Leung showed that if $(e_{i})_{i}$ is the canonical basis for $\ell_{p}$, $2\leq p<\infty$, then $J(e_{i})^{\prime}$, the dual of James space $J(e_{i})$, is symmetrically Arens regular but not Arens regular. Observe that $J(e_{i})$ is not a Banach lattice since it is a non-reflexive space containing no copies of $c_{0}$ and $\ell_{1}$ (see [1, Theorem 4.71]).
Now, we prove a lemma that will be needed later in this section.
Lemma 2.2.
@@
Proof.
Let $A\colon E\times E\to\mathbb{R}$ be a positive symmetric bilinear form. On the one hand, since $A$ is symmetric, $A=A^{t}$. On the other hand, as $E^{*}$ has order continuous norm, it follows from Proposition 2.1 that $\overline{A}=A^{t***t}$. Thus, for all $x^{**},y^{**}\in E^{**}$
$\overline{A}(x^{**},y^{**})=A^{t***t}(x^{**},y^{**})=A^{t***}(y^{**},x^{**})=A^{***}(y^{**},x^{**})=\overline{A}(y^{**},x^{**}),$
proving that $\overline{A}$ is symmetric for every positive symmetric bilinear form $A\colon E\times E\to\mathbb{R}$. Moreover, the variables in a positive symmetric $n$-linear form $A\colon E^{n}\to\mathbb{R}$ can be permuted in pairs in the extensions (see, e.g., [31, p. 52]) to get that $\overline{A}$ is symmetric.
Now, if $A\colon E^{n}\to F$ is a positive symmetric $n$-linear operator, then, for every $0\leq y^{\ast}\in F^{\ast}$, $y^{\ast}\circ A\colon E^{n}\to\mathbb{R}$ is a positive symmetric $n$-linear form, and by the above case $J_{F^{\ast}}(y^{\ast})\circ\overline{A}=\overline{y^{\ast}\circ A}$ is also a positive symmetric $n$-linear form. In particular, for every $0\leq y^{\ast\ast\ast}\in F^{\ast\ast\ast}$, we get that $0\leq y^{\ast\ast\ast}|_{F^{\ast}}\in F^{\ast}$, and hence $y^{\ast\ast\ast}\circ\overline{A}$ is symmetric. Finally, let $x_{1}^{**},\dots,x_{n}^{**}\in E^{**}$ and $\sigma\in S_{n}$. If $\overline{A}(x_{1}^{**},\dots,x_{n}^{**})\neq\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**})$, then there exists $y^{***}\in F^{***}$ such that
$y^{***}\circ\overline{A}(x_{1}^{**},\dots,x_{n}^{**})$ $=y^{***}(\overline{A}(x_{1}^{**},\dots,x_{n}^{**}))$
$\neq y^{***}(\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**}))=y^{***}\circ\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**}),$
a contradiction, proving that $\overline{A}$ is symmetric.∎
@@
The following lemma will be necessary in our discussion. Since we did not find it anywhere in the literature, we will include its proof for convenience.
Lemma 2.3.
If $A\colon E_{1}\times\cdots\times E_{n}\to F$ is a positive $n$-linear map, then its Arens extension $\overline{A}\colon E_{1}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\to F^{\ast\ast}$ is also positive.
Proof.
Let us recall the construction of $\overline{A}$. Let $A^{\ast}\colon F^{\ast}\times E_{1}\times\cdots\times E_{n-1}\to E_{n}^{\ast}$ be defined by
$A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1})(x_{n})=y^{\ast}(A(x_{1},\dots,x_{n})).$
Then $A^{\ast\ast}\colon E_{n}^{\ast\ast}\times F^{\ast}\times E_{1}\times\cdots\times E_{n-2}\to E_{n-1}^{\ast}$ is given by
$A^{\ast\ast}(x_{n}^{\ast\ast},y^{\ast},x_{1},\dots,x_{n-2})(x_{n-1})=x_{n}^{\ast\ast}(A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1}))$
Supposing that $A^{\ast(n)}\colon E_{2}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\times F^{\ast}\to E_{1}^{\ast}$ is defined inductively, the Arens $\overline{A}=A^{\ast(n+1)}\colon E_{1}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\to F^{\ast\ast}$ extension of $A$ is given by
$A^{*(n+1)}(x_{1}^{\ast\ast},\dots,x_{n}^{\ast\ast})(y^{\ast})=x_{1}^{\ast\ast}(A^{*(n)}(x_{2}^{\ast\ast},\dots,x_{n}^{\ast\ast},y^{\ast})).$
Note that for all $y^{\ast}\in(F^{\ast})^{+}$ and $x_{i}\in E_{i}^{+}$, we get that
$A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1})(x_{n})=y^{\ast}(A(x_{1},\dots,x_{n}))\geq 0,$
which implies that $A^{\ast}\geq 0$. By the same reasoning, $A^{\ast\ast}\geq 0$. Inductively, we get that $\overline{A}\geq 0$.∎
@@
We begin by proving that $\overline{\varphi}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. To see this, let $P\colon E\to G$ be a positive $n$-homogeneous polynomial. We will prove that $\tilde{\varphi(T_{P})}(y_{1},\dots,y_{n})\geq 0$ in $G^{\ast\ast}$ for all $y_{1},\dots,y_{n}\in F^{+}$ Indeed, given $y_{1},\dots,y_{n}\in F^{+}$ and $0\leq z^{\ast}\in G^{\ast}$, we have that
@@
because $\overline{z^{\ast}\circ T_{P}}\geq 0$ (due to $z^{\ast}\circ T_{P}\geq 0$), $\varphi^{\ast}\geq 0$ (due to $\varphi\geq 0$) and $J_{F}\geq 0$ (see [26, Proposition 1.4.5(ii)]), proving that $\tilde{\varphi}(T_{P})\colon F^{n}\to G^{\ast\ast}$ is a positive $n$-linear map. Consequently, $\tilde{\varphi}(T_{P})^{s}$ is also a positive $n$-linear map (see [27, Proposition 1.6]). Since $\tilde{\varphi}(T_{P})^{s}$ is the symmetric $n$-linear form associated to $\overline{\varphi}(P)$, we conclude that $\overline{\varphi}(P)$ is a positive $n$-homogeneous polynomial. Therefore, $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is positive. Since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we also have that $\overline{\varphi^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{**})\to\mathcal{P}^{r}(^{n}E;G^{****})$ is also a positive operator.
Finally, since $E^{*}$ has order continuous norm, we get by Lemma 2.2 that $\overline{T_{P}}$ is symmetric for every positive $n$-homogeneous polynomial $P\colon E\to G$. Thus, [14, Lemma 1.1(b)] yields that $\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=P$ holds for every $P\in\mathcal{P}^{+}(^{n}E;G)$, and hence
@@
The first problem in Theorem 2.4 is that $\overline{\varphi}(P)$ may not take its values in $G$ (see [14, p. 285]). The second one is that we cannot ensure that the image of $\overline{\varphi}$ is a Banach lattice. However, when we consider polynomials taking values on a dual Banach lattice, we obtain the following.

### v2

@@ Abstract
We study Díaz-Dineen’s problem for regular homogeneous vector-valued polynomials. In particular, we prove that if $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We also study the analogous problem for the classes of regular compact, regular weakly compact, orthogonally additive and regular nuclear polynomials.
@@ 1 Introduction and background
In [18], J. C. Díaz and S. Dineen posed the following question: If $X$ and $Y$ are dual isomorphic Banach spaces, that is $X^{\ast}$ and $Y^{\ast}$ are isomorphic as Banach spaces, are $\mathcal{P}(^{n}X)$ and $\mathcal{P}(^{n}Y)$ isomorphic as Banach spaces? Although it is not known if the above question has a positive answer, some partial answers were presented in [16, 18, 29]. In particular, J. C. Díaz and S. Dineen proved that the problem has a positive solution whenever $X^{\ast}$ has both the Schur property and the approximation property (see [18, Proposition 4]). In [16], F. Cabello Sánchez, J. Castillo and R. García generalized Díaz-Dineen results by providing a positive solution for this problem whenever $X$ is Arens-regular or whenever $X$ and $Y$ are stable. The same result was proven, independently, by S. Lassalle and I. Zalduendo in [29] who also addressed the same question concerning other classes of polynomials. Further, in [17], D. Carando and S. Lassalle extended these results to spaces of vector-valued homogeneous polynomials.
@@
Motivated by the increasing interest in studying homogeneous polynomials on Banach lattices (see, e.g., [6, 7, 8, 9, 10, 11, 13, 24, 31]), the following question arises naturally: Given two dual isomorphic Banach lattices $E$ and $F$, that is $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic, under which conditions are $\mathcal{P}^{r}(^{n}E)$ and $\mathcal{P}^{r}(^{n}F)$ lattice isomorphic for every $n\in\mathbb{N}$? By adapting the proofs from [17] and [29], we were able to provide a positive answer whenever $E^{*}$ and $F^{*}$ have order continuous norms. In particular, we prove that if $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We observe that, in our setting, the assumption of Arens regularity is replaced with the requirement that the norm on the dual is order continuous. This replacement is justified by the fact that a positive version of Arens regularity for $E$ is equivalent to the order continuity of the norm on $E^{*}$. Furthermore, as we will demonstrate in Proposition 2.1, positive Arens regularity coincides with classical Arens regularity and with positive symmetric Arens regularity, highlighting a noteworthy contrast with the Banach space case.
@@
As an application, we obtain conditions so that $\mathcal{P}_{\cal K}^{r}(^{n}E;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}E;G^{*})$) and $\mathcal{P}_{\cal K}^{r}(^{n}F;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}F;G^{*})$) are isomorphic, where $\mathcal{P}_{\cal K}^{r}$ and $\mathcal{P}_{\cal W}^{r}$ denote, respectively, the classes of regular compact and regular weakly compact polynomials. We also consider the class of orthogonally additive polynomials. In Section 3, we investigate this problem for a new class of polynomials. Specifically, inspired by ideas from [4], we define the class of regular nuclear $n$-homogeneous polynomials, denoted by $\mathcal{P}_{\cal N}^{r}$, and prove that if $G$ is Dedekind complete and has the lattice approximation property (LAP), and $E^{*}$ and $F^{*}$ are lattice isomorphic both having the LAP, then $\mathcal{P}_{\cal N}^{r}(^{n}E;G)$ and $\mathcal{P}_{\cal N}^{r}(^{n}F;G)$ are lattice isomorphic for every $n\in\mathbb{N}$.
@@
Before proceeding, we introduce some definitions and notation. Throughout this paper, $X$ and $Y$ denote Banach spaces, $E$ and $F$ denote Banach lattices, $E^{+}$, $B_{E}$ and $S_{E}$ denote, respectively, the positive cone, the closed unit ball, and the unit sphere of $E$. Given Banach spaces $X_{1},\dots,X_{n}$ and $Y$, the Banach space of all continuous $n$-linear operators $A\colon X_{1}\times\cdots\times X_{n}\to Y$ is denoted by $\mathcal{L}(X_{1},\dots,X_{n};Y)$. In the case that $F=\mathbb{K}$, we simply write $\mathcal{L}(X_{1},\dots,X_{n})$. Given Banach lattices $E_{1},\dots,E_{n}$ and $F$, a $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$ is said to be positive if $A(x_{1},\dots,x_{n})\geq 0$ for all $x_{1}\in E_{1}^{+},\dots,x_{n}\in E_{n}^{+}$. The difference of two positive $n$-linear operators is called a regular $n$-linear operator, and the set of all regular $n$-linear operators from $E_{1}\times\cdots\times E_{n}$ into $F$ is denoted by $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$. Whenever $F$ is Dedekind complete, $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$ is a Banach lattice with the regular norm $\|A\|_{r}=\||A|\|$, where $|A|$ denotes the absolute value of the regular $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$. The positive projective tensor product, or the Fremlin tensor product of $E_{1},\dots,E_{n}$, denoted by $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$, is the Banach lattice generated by considering the completion of the vector lattice tensor product $E_{1}\overline{\otimes}\cdots\overline{\otimes}\,E_{n}$ with respect the positive projective tensor norm:
@@
is an isometric isomorphism and a lattice hommomorphism (see [14, Proposition 3.3]).
An $n$-homogeneous polynomial between Riesz spaces $P\colon E\to F$ is positive if its associated symmetric multilinear operator $T_{P}\colon E^{n}\to F$ is positive. The difference of two positive $n$-homogeneous polynomials is called a regular homogeneous polynomial, and the set of such polynomials is denoted by $\mathcal{P}^{r}(^{n}E,F)$. When $F$ is the scalar field we simply write $\mathcal{P}^{r}(^{n}E)$. If $E$ and $F$ are Banach lattices with $F$ Dedekind complete, then $\mathcal{P}^{r}(^{n}E,F)$ is a Banach lattice with the regular norm $\|P\|_{r}=\||P|\|$, where $|P|$ denotes the absolute value of the regular $n$-homogeneous polynomial $P\colon E\to F$. The $n$-fold positive projective symmetric tensor product of $E$, denoted by $\widehat{\otimes}_{n,s,|\pi|}E$, is a Banach lattice endowed with the positive projective symmetric tensor norm:
@@
where $\otimes^{n}x=x\,\otimes\stackrel{{\scriptstyle n}}{{\cdots}}\otimes\,x$ for every $x\in E$. Let $\theta_{n}\colon E\to\widehat{\otimes}_{n,s,|\pi|}E$ be the canonical $n$-homogeneous polynomial given by $\theta_{n}(x)=\otimes^{n}x$. We note that $\theta_{n}$ is a lattice homomorphism. For every $P\in\mathcal{P}^{r}(^{n}E,F)$ there exists a unique regular linear operator $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$, called the linearization of $P$, such that $P(x)=P^{\otimes}(\theta_{n}(x))$ for every $x\in E$. The operator
@@
is an isometric isomorphism and a lattice homomorphism [14, Proposition 3.4].
We refer the reader to [1, 33] for background on Banach lattices, to [21] for Banach space theory, to [19] for polynomials on Banach spaces, to [14, 32] for positive/regular multilinear operators or polynomials, and to [14, 22, 23] for tensor products between vector lattices.
@@ 2 Díaz-Dineen problem for regular polynomials
This process is repeated until we obtain $\overline{A}:=A^{(n+1)*}\colon E_{1}^{**}\times\cdots\times E_{n}^{**}\to F^{**}$, the $(n+1)$-Arens adjoint of $A$ which is called the Arens-extension of $A$ (see [9, Section 2] for more details). For the remainder of the paper, we shall call this extension the Arens extension, bearing in mind that it coincides with its Aron-Berner extension (see [3]). Whenever $X=Y$, the transpose of $A$ is the bilinear map $A^{t}\colon X\times X\to Z$ given by $A^{t}(x,y)=A(y,x)$ for all $x,y\in X$. Still under the assumption that $X=Y$, AA is said to be Arens-regular (or simply regular) if $A^{t***t}=A^{***}$. It is important to note that we have two different concepts of regularity for multilinear operators, one in the lattice sense presented in the Introduction and the one in the extension sense. It follows from [37, Theorem 2.2] that a bilinear form $A\colon X\times X\to\mathbb{R}$ is Arens-regular if and only if the bounded linear operator $S_{A}\colon X\to X^{*}$ given by $S_{A}(x)(y)=A(x,y)$ is weakly compact.
Recall that a Banach space $X$ is called Arens-regular if all linear operators from $X$ into $X^{*}$ are weakly compact. To our knowledge, the first appearance of a positive version of Arens-regularity was in [15], when G. Buskes and R. Page gave necessary and sufficient conditions so that every positive bilinear form $E\times E\to\mathbb{R}$ is Arens-regular. The following result collects the characterizations obtained so far by G. Buskes, R. Page, G. Emmanuele, E. Saab and P. Saab, and proves that positive Arens-regularity does not give rise to a genuinely new concept as it is equivalent to classical Arens-regularity.
@@
Since $B$ is positive, we may consider the positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ given by $A(x,y)=B(\Pi(x),\Pi(y))$ for all $x,y\in E$. Let us see that $A$ cannot be Arens-regular, showing that its associated operator $S_{A}\colon E\to E^{*}$ is not weakly compact: as $\{S_{B}(e_{2n}):n\in\mathbb{N}\}$ is not a weakly precompact subset of $\ell_{1}^{*}=\ell_{\infty}$ (see [3, p. 83]), there exists $\varphi\in\ell_{\infty}$ such that $(\varphi(S_{B}(e_{2n})))_{n}$ does not have any Cauchy subsequence. If $x_{n}\in E$ is such that $\Pi(x_{n})=e_{2n}$, then $(\varphi(S_{B}(\Pi(x_{n}))))_{n}$ does not have any Cauchy subsequence. By extending $\varphi\colon\ell_{1}\to\mathbb{R}$ to $E$ in such a way that it vanishes on the complement, we obtain that $\{\Pi^{*}\circ S_{B}\circ\Pi(x_{n}):n\in\mathbb{N}\}$ is not weakly precompact in $E^{*}$, and since $S_{A}=\Pi^{*}\circ S_{B}\circ\Pi$, we conclude that $S_{A}$ is not weakly compact.
[…]
It is worth mentioning that Proposition 2.1 shows that the notions of Arens-regularity, positive Arens-regularity, and positive symmetric Arens-regularity are all equivalent in the Banach lattice setting. This result highlights an important difference between the theory of Banach spaces and the theory of Banach lattices. In [30], D. Leung showed that if $(e_{i})_{i}$ is the canonical basis for $\ell_{p}$, $2\leq p<\infty$, then $J(e_{i})^{\prime}$, the dual of James space $J(e_{i})$, is symmetrically Arens regular but not Arens regular. Observe that $J(e_{i})$ is not a Banach lattice since it is a non-reflexive space containing no copies of $c_{0}$ or $\ell_{1}$ (see [1, Theorem 4.71]).
Our next result is an immediate consequence of Proposition 2.1 and [39, p. 52].
Corollary 2.2.
@@
In the proof of our next results, we will use the fact that the Arens extension of every positive $n$-linear map is also positive [5, Theorem 2.2(c)]. A proof of this result can be found in the doctoral dissertation of L. Garcia [26, Theorem 2.1.3].
@@
We begin by proving that $\overline{\varphi}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. To see this, let $P\colon E\to G$ be a positive $n$-homogeneous polynomial. We will prove that $\tilde{\varphi}(T_{P})(y_{1},\dots,y_{n})\geq 0$ in $G^{\ast\ast}$ for all $y_{1},\dots,y_{n}\in F^{+}$. Indeed, given $y_{1},\dots,y_{n}\in F^{+}$ and $0\leq z^{\ast}\in G^{\ast}$, we have that
@@
because $\overline{z^{\ast}\circ T_{P}}\geq 0$ (due to $z^{\ast}\circ T_{P}\geq 0$), $\varphi^{\ast}\geq 0$ (due to $\varphi\geq 0$) and $J_{F}\geq 0$ (see [33, Proposition 1.4.5(ii)]). This proves that $\tilde{\varphi}(T_{P})\colon F^{n}\to G^{\ast\ast}$ is a positive $n$-linear map, and consequently $\tilde{\varphi}(T_{P})^{s}$ is also a positive $n$-linear map (see [34, Proposition 1.6]). Since $\tilde{\varphi}(T_{P})^{s}$ is the symmetric $n$-linear form associated to $\overline{\varphi}(P)$, we conclude that $\overline{\varphi}(P)$ is a positive $n$-homogeneous polynomial. Therefore, $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is positive. Since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we also have that $\overline{\varphi^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{**})\to\mathcal{P}^{r}(^{n}E;G^{****})$ is also a positive operator.
Finally, since $E^{*}$ has order continuous norm, we get by Corollary 2.2 that $\overline{T_{P}}$ is symmetric for every positive $n$-homogeneous polynomial $P\colon E\to G$. Thus, [17, Lemma 1.1(b)] yields that $\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=P$ holds for every $P\in\mathcal{P}^{+}(^{n}E;G)$, and hence
@@
The first limitation of Theorem 2.3 is that $\overline{\varphi}(P)$ may not take its values in $G$ (see [17, p. 285]). The second is that we cannot ensure that the image of $\overline{\varphi}$ is a Banach lattice. However, when we consider polynomials taking values on a dual Banach lattice, we obtain the following.